OpenStudy (anonymous):

Point A, located at (-8, 3), is the location of the storage closet and point C, located at (12, -4), is the director's office. Find point B, the location of the art supplies, if it is 2 over 5 the distance from point A to point C.

OpenStudy (anonymous):

I DONT GET IT

OpenStudy (jdoe0001):

so... where do you think we'd find our "art supplies" section? we know is between point A and C what's the distance between A and C anyway?

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ A&({\color{red}{ -8}}\quad ,&{\color{blue}{ 3}})\quad C&({\color{red}{ 12}}\quad ,&{\color{blue}{ -4}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\) point B seems to be in between A and C, from A , B seems to be \(\bf \cfrac{2}{5}th\) from that distance

OpenStudy (anonymous):

Thats where I get lost, Im not so sure how to find the distance between the two

OpenStudy (anonymous):

Oh wait, yes I am sorry hold on

OpenStudy (anonymous):

is it 21.2

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

are you doing geometry?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

2/5 of 21.2 will be a distance but rereading you.... seems it's asking for the actual x,y coordinates for point B

OpenStudy (anonymous):

yeah it says the middle of A and C

OpenStudy (jdoe0001):

well, it's not the middle, is 2/5 the way from A to C

OpenStudy (jdoe0001):

so is a matter of 2 ratios

OpenStudy (jdoe0001):

if I read you right

OpenStudy (anonymous):

oh yeah okay sorry my brain is fried ive been spending 2 hours on this project... how do I know where 2/5 of the way is?

OpenStudy (jdoe0001):

lemme show... you one... from 2 days... same thing, 2 ratios.... see if you can follow the example -> http://openstudy.com/study#/updates/52ec21c6e4b0a362a8947591

OpenStudy (jdoe0001):

that one happens to use 1:3 ratios, yours is 2:5

OpenStudy (anonymous):

I looked at it, still dont get it.

OpenStudy (jdoe0001):

ok.... lemme rewrite it for this one

OpenStudy (anonymous):

thanks

OpenStudy (jdoe0001):

 \(\bf \color{blue}{ A(-8,3)\qquad C(12,-4)\\ \quad \\ \quad \\ \cfrac{AB}{CB}=\cfrac{ratio1}{ratio2}\implies ratio2\cdot AB=ratio1\cdot CB\quad \textit{dividing by B}\\ \quad \\ ratio2\cdot A=ratio1\cdot C}\implies 5(-8,3)=2(12,-4)\\ \quad \\\qquad B=\left(\cfrac{\textit{sum of "x" values}}{ratio1+ratio2}\quad ,\quad \cfrac{\textit{sum of "y" values}}{ratio1+ratio2}\right)\\ \quad \\ \qquad thus\qquad \\ \quad \\ B=\left(\cfrac{(5\cdot -8)+(2\cdot 12)}{2+5}\quad ,\quad \cfrac{(5\cdot 3)+(2\cdot -4)}{2+5}\right)\)

OpenStudy (anonymous):

the asnwer is -0.9?

OpenStudy (jdoe0001):

the answer would be an x,y coordinate set

OpenStudy (anonymous):

oh no, uh I got -1.6 for this first and 7 for the second so im doing something wrong